OpenStudy (anonymous):

A .24 kg volleyball approaches Tina with a velocity of 3.8 m/s. Tina bumps the ball, giving it a speed of 2.4 m/s but in the opposite direction. What average force did she apply if the interaction time between her hands and the ball was 0.025 s?

OpenStudy (espex):

\[F_{avg}=ma_{avg} \rightarrow m\frac{\Delta v}{\Delta t}\] \[.24kg \times \frac{-2.4\frac{m}{s}-3.8\frac{m}{s}}{0.025-0}=-59.52N\] The negative sign just means that it changed direction.